40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 186 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  7599   Tue Oct 23 17:30:33 2012 jamie, nic, jenne, raji, manasaUpdateAlignmentInitial attempts to fix IFO alignment

We went into the vertex today to see about fixing the alignment.  The in-air access connector is in place, and we took heavy doors off of BS, ITMY, and ETMY chambers.

We started by looking at the pointing from the PZTs.  Manasa and Raji hooked up HV power supplies to the PZTs and set them to the middle of their ranges (75 V).

We installed a target on the BS cage, and new "free standing" targets made special by Steve for the SOSs on ITMY and ETMY.

Using a free-standing aperture target we looked at the beam height before PZT2.  It was a little high, so we adjusted it with PZT1.  Once that was done we looked at the beam height at PR2, and adjusted that height with PZT1.

We then tried to use the hysteresis in PR2 to adjust the beam height at ITMY.  Pushing just a little bit at the top or bottom of PR2 would repoint the beam in pitch.  This sort of works, but it's stupid.  Using this method we got the beam more or less centered vertically at ITMY.

We moved on to ETMY with the idea that we would again use the hysteresis in PR3 to get the vertical pointing to the ETM correct.  This was a good demonstration of just how stupid the tip-tilts really are.  Just touching slightly at the top or bottom or PR3 we could completely change the pointing at ETMY, by mili-radians (~4 cm over 40m).

At this point I cried foul.  This is not an acceptable situation.  Very little stimulation to the tip-tilts can repoint the beam inside the PR cavity.

Steve says that the TT weights, which will attach to the base of the TT mirror mounts and should help keep the mirrors vertical and not hysteretic, are being baked now and should be available tomorrow.  We therefore decided to stop what we were doing today, since we'll have to just redo it all again tomorrow once the weights are installed.

 

  7598   Tue Oct 23 17:12:30 2012 DenUpdateSUS PITCH damping needed

 Koji and Steve pointed out that previous design  of a damping bracket was a bit complicated to manufacture. So I made it simpler and also added a tap hole for original yaw damping. We'll give drawing to Mike in the machine shop tomorrow morning.

I've purchased K&J magnets for eddy current damping, they should be here in 2 days. 

Attachment 1: simple_drawing.PDF
simple_drawing.PDF
  7597   Tue Oct 23 16:18:03 2012 SteveUpdateVACRoughing pumps on for oil change

Leybold D30AQuote:

PR1, PR2 and RP3 turned on for warming up for oil change. Oil changed with 3.2L of MVT-19  fluid in each. This substitute for HE-175 will be used next time.

To finish this job tomorrow: 1, check oil creeping upstream  2, change air filter of air purge if pressure drops <350 mTorr  3, measure venting time of pump

 . Leybold D30A manual is here. Exhaust filter traps were drained. No oil creeping was found. The pump venting time is < 1 minute.

PR-2 needs new secu-valve. PR1 & 2 are in excellent condition.

Attachment 1: oilchange.jpg
oilchange.jpg
  7596   Tue Oct 23 10:24:42 2012 AyakaUpdateAdaptive Filteringmicrophone noise

Quote:

Quote:

I will do some experiments on acoustic noise canceling during my stay.
Now I am planning to cancel acoustic noise from PMC and see how the acoustic noise work and how we should place microphones.a

First, I measured the noise in microphones and its circuit.
mic_noise2.png
-blue, green, red, solid lines; microphone signals
-blue, green, red, dashed lines; un-coherent noise in signals
-yellow, black, solid lines; circuit noise (signal input is open, not connected to the microphones)

We can see the acoustic signal above 1 Hz, and the circuit does not seem to limit its sensitivity. But I do not know why yellow and black is so different. I will check it tomorrow.

Hi, Ayaka.  It would be good if you could give a little bit more detail about this plot:

  • What exactly are the "signals"?  Are you making a sound somehow?  If so, what is producing the sound?  What is it's spectrum?
  • Are the blue/green/red traces from three different microphones?
  • Coherence usually implies a comparison between two signals.  Is something being compared in the dashed traces?
  • Are the yellow and black traces from different amplifiers?
  • What are the units of the Y axis?

 

 Sorry for my poor explanation.

I measured this by the same way as you measured the instrumental noise of seismometers.
I put the three microphones at the same place so that the three can hear the same sound. I did not make any sounds, just put them in the lab.
The signals from microphones are all amplified by the circuit.
And I took the correlations of each signals and two others and got the noise (dashed lines) by subtracting the correlated signal from the original signal.

So,
-The signal is the acoustic sound in the lab, amplified by the circuit.
-Three lines are from three different microphones.
-Dashed lines are subtraction of coherent signal from the original.
-Yellow and black lines are from different amplifiers in the same circuit box. The circuit has 6 channels.
-I did not calibrate the signals I got by DTT since I do not know the calibration factor now. It is just the number I got from the real time system.

 

  7595   Tue Oct 23 09:37:33 2012 SteveUpdateSUS passive TT adjustment correction

Quote:

Tip Tilt pitch adjustment on existing-in vacuum suspension. This can be added by a simple installation of a 1.25" long 2-56 threaded rod with nuts.

 We should check that  their sus wire diameter are 0.0017"  All 2-56 hardware are in and  Bob is cleaning them.

Attachment 1: IMG_1745.JPG
IMG_1745.JPG
  7594   Tue Oct 23 08:45:00 2012 SteveUpdateVACvent day 3

Quote:

Quote:

Quote:

Quote:

 

 PSL shutter closed, manual block in place, HV turned off. P1 is at 200 Torr now.  Jenne is taking over here.

 Valves closed, 500 torr.  Steve will finish off Monday morning, then we'll take off doors and get to work.

 We are almost at atm.  P1 750 Torr,  We are slowly reaching equilibrium.

 The access connector was removed, BS and ITMY chambers were opened yesterday.

New one piece aluminum shell access connector installed. Jenne lowered the PSL out put into MC to ~100mW

Attachment 1: ventd4.png
ventd4.png
Attachment 2: IMG_1743.JPG
IMG_1743.JPG
  7593   Tue Oct 23 01:46:53 2012 DenUpdateSUS PITCH damping needed

Quote:

 

Too bad - I thought it would at least give a little damping. Since we want the viscous-like energy loss to be ~49x larger, we need to have the field modulation in the damper (not dumper) increase by ~7.

 I've made SolidWorks models of damping bracket and eddy current disk. They will me manufactured and used instead of old ones. New bracket will be mounted in exactly the same place where the old one was. Drawings might not be complete but all dimensions are in the models so we can fix drawing tomorrow before going to machine shop.

I think we can use ring magnets for passive damping. Then we won't have the vent problem. I've found some at K&J Magnetics, we can get them any time. Magnets are Ni-Cu-Ni (fine for vacuum?) Diameter is 3/8'' with advertised tolerence 0.004'', so they should fit the holes.

Attachment 1: Mirror_Holder_ECD_NEW_DRAWING.PDF
Mirror_Holder_ECD_NEW_DRAWING.PDF
Attachment 2: TT_DAMPING_DRAWING.PDF
TT_DAMPING_DRAWING.PDF
  7592   Tue Oct 23 00:51:41 2012 JamieUpdateAdaptive Filteringmicrophone noise

Quote:

I will do some experiments on acoustic noise canceling during my stay.
Now I am planning to cancel acoustic noise from PMC and see how the acoustic noise work and how we should place microphones.a

First, I measured the noise in microphones and its circuit.
mic_noise2.png
-blue, green, red, solid lines; microphone signals
-blue, green, red, dashed lines; un-coherent noise in signals
-yellow, black, solid lines; circuit noise (signal input is open, not connected to the microphones)

We can see the acoustic signal above 1 Hz, and the circuit does not seem to limit its sensitivity. But I do not know why yellow and black is so different. I will check it tomorrow.

Hi, Ayaka.  It would be good if you could give a little bit more detail about this plot:

  • What exactly are the "signals"?  Are you making a sound somehow?  If so, what is producing the sound?  What is it's spectrum?
  • Are the blue/green/red traces from three different microphones?
  • Coherence usually implies a comparison between two signals.  Is something being compared in the dashed traces?
  • Are the yellow and black traces from different amplifiers?
  • What are the units of the Y axis?

 

  7591   Tue Oct 23 00:41:55 2012 JenneUpdateAlignmentMC locked, spots centered

[Jenne, Raji]

We replaced the MC Refl path BS with the Y1, as usual, so that the full ~100mW goes to the REFL PD, so we don't have WFS or MC refl camera. 

The MC spots were all outside of 1mm, and some were beyond 2mm (for MC1,3, P,Y....MC2 is of course free since we have more DoFs than we need), so we touched (very, very slightly) the zigzag mirrors on the PSL table.  We realigned the MC, and now the spots are centered to my satisfaction.

MC1,2,3 Pit,    MC1,2,3 Yaw (in mm):

[0.46444020918749457, 8.2634316545130009, -0.41417975237831089, -0.89401481457980592, -0.9323196976382162, -1.543145765853893]

MC2 is way off in pitch, according to this measurement, and it's been consistently going down as we move the MC2 spot in the same direction (up on the monitor), but since we started at +15mm and are now at +8, and we've gone quite a ways, I'm not sure that we really want to go all the way to 0.  Anyhow, MC1 and MC3 are the ones which define our input pointing, so we're quitting for tonight.

We will turn on the PZTs and begin with the official vent list for dummies tomorrow.

Attachment 1: Screenshot-Figure_1.png
Screenshot-Figure_1.png
  7590   Mon Oct 22 21:20:36 2012 JenneUpdatePSLBeam attenuation optics in place

[Jenne, Raji (before dinner)]

We put the beam attenuation optics in place.  Before putting any optics down, I centered the IOO QPDs, then adjusted the HWPs and PBS such that we remained centered on those QPDs.

Now, I'm about to unblock the beam and let ~100mW into the vacuum so I can lock the MC.  Steve and Manasa were putting on the light access connector when I left earlier, so I'm excited to use it!

  7589   Mon Oct 22 20:44:49 2012 AyakaUpdateAdaptive Filteringmicrophone noise

I will do some experiments on acoustic noise canceling during my stay.
Now I am planning to cancel acoustic noise from PMC and see how the acoustic noise work and how we should place microphones.

First, I measured the noise in microphones and its circuit.
mic_noise2.png
-blue, green, red, solid lines; microphone signals
-blue, green, red, dashed lines; un-coherent noise in signals
-yellow, black, solid lines; circuit noise (signal input is open, not connected to the microphones)

We can see the acoustic signal above 1 Hz, and the circuit does not seem to limit its sensitivity. But I do not know why yellow and black is so different. I will check it tomorrow.

  7588   Mon Oct 22 11:26:12 2012 ranaUpdateSUS PITCH damping needed

 

Too bad - I thought it would at least give a little damping. Since we want the viscous-like energy loss to be ~49x larger, we need to have the field modulation in the damper (not dumper) increase by ~7.

  7587   Mon Oct 22 09:10:07 2012 SteveUpdateVAC@750Torr

Quote:

Quote:

Quote:

 

 PSL shutter closed, manual block in place, HV turned off. P1 is at 200 Torr now.  Jenne is taking over here.

 Valves closed, 500 torr.  Steve will finish off Monday morning, then we'll take off doors and get to work.

 We are almost at atm.  P1 750 Torr,  We are slowly reaching equilibrium.

Attachment 1: ventOct222012.png
ventOct222012.png
  7586   Sat Oct 20 20:37:55 2012 DenUpdateSUS PITCH damping needed

Quote:

 

 We need stronger magnets if we want eddy current dumping down to Q~1.

 I've inserted 10mm * 10mm magnets to the 4 corner holes on the front side of the mirror frame according to actuation magnets polarity. I realigned TT and measured Q factor for pitch and yaw, it was 5-10.

DSC_4778.JPG     pitch.JPG   yaw.JPG

 

I was able to do it for 1 TT only, because others have smaller (~0.1 mm) hole diameter and magnets can't go inside. I tried to warm holes up to 850 F but still was not able to insert a magnet.

DSC_4776.JPG

  7585   Sat Oct 20 01:23:56 2012 DenUpdateSUS PITCH damping needed

Quote:

Measurements confirm that damping using small actuation magnets reduces pitch Q by a factor of 4 and is not enough.

 I've tested the idea to use coils as eddy current dampers. I terminated them with a wire and measured Q factor during the ringdown test. Sadly, I did not see any significant damping and Q was ~150. We need stronger magnets if we want eddy current dumping down to Q~1.

P1010059.jpg      P1010060.jpg    P1010058.jpg

  7584   Fri Oct 19 19:26:42 2012 JenneUpdateVACvent has started

Quote:

Quote:

 

 PSL shutter closed, manual block in place, HV turned off. P1 is at 200 Torr now.  Jenne is taking over here.

 Valves closed, 500 torr.  Steve will finish off Monday morning, then we'll take off doors and get to work.

  7583   Fri Oct 19 18:30:07 2012 DenUpdateSUS PITCH damping needed

CQuote:

 

 From this data I've estimated Q factor for yaw (135 vs 88) and pitch (192 vs 77) (original vs added pitch damping).

 I've made a more precise measurement of pitch damping using spectrum analyzer.

damp.png   pitch.png

Measurements confirm that damping using small actuation magnets reduces pitch Q by a factor of 4 and is not enough.

  7582   Fri Oct 19 16:45:39 2012 SteveUpdateVACvent has started

Quote:

 

 VENT NOW and FIX ALIGNMENT!

 PSL shutter closed, manual block in place, HV turned off. P1 is at 200 Torr now.  Jenne is taking over here.

  7581   Fri Oct 19 16:24:39 2012 ranaUpdateLockingAlignment is really bad??

 

 VENT NOW and FIX ALIGNMENT!

  7580   Fri Oct 19 12:45:12 2012 DenUpdateCDSc1lsc is up after reboot
  7579   Fri Oct 19 01:21:42 2012 EvanUpdateLockingAligning PZTs, PRM

[Evan, Jenne]

Tonight we made an attempt at getting the PRM + ITMY aligned with correct input pointing. We steered the good PZT so that the input beam makes it through the aperture in front of ETMY. We then aligned the PRM so that the retroreflection of the input beam makes it back into the Faraday. After that we tried dithering the alignment of ITMY and the beamsplitter to see if we could see a spot flash across the AS port, but we saw nothing.

For the PRM alignment we set up a camera looking into the window at the Faraday in the IOO chamber; it's called FI_BACK. We stole a 50mm lens from the ETMY face camera.

We also tried looking for beam on IP_POS and IP_ANG. When the input beam is aligned to pass through the ETMY aperture, we can see beam on the steering mirrors preceding IP_POS, but it hits a mirror mount. When the input beam is aligned as it was on Monday, it clips on the ETMY aperture but makes it further along the IP_POS optical path.   In both cases, we weren't able to see any beam coming out for IP ANG. 

  7578   Fri Oct 19 01:11:18 2012 JenneUpdateIOOMCL, WFS triggers

Quote:

Quote:

 

 I've redone the WFS triggers.  I left the MCL trigger alone (for now....I'll come back to it). 

The trigger was setup such that (a) it was totally unclear what was going on, by looking at the WFS screen.  Koji and I spent some time confused before I remembered that Den did this work recently.  Also, for some reason, the triggers were just plain thresholding, not a schmidt trigger, so any time the cavity flashed, the WFS came on.  Since the cavity can flash before the mcdown script has a chance to turn off the WFS servos, the outputs of the WFS filters are trying to output thousands of counts, and the signal goes through any time the cavity flashes.  Not so good.

 Your schmitt trigger has 2 threshold values - min and max. Set thresholding value in my trigger to the max of your schmitt trigger and you get the same behavior for MC,  triggers are not supposed to turn anything on in this realization as they do for locking with flashing.

 The problem is that the WFS were being engaged with your triggers every time the MC flashed.  That wasn't a schmidt trigger thing, but I like the schmidt trigger better anyway.

 

Anyhow, it's turned on, and it works really well.  It's kind of awesome.  I'm really excited to start using the wait block to start pushing even more of the locking out of scripts and into the real time system.

  7577   Fri Oct 19 00:55:35 2012 JenneUpdateComputersc1lsc is down (at least all of the models)

When Evan and I were dithering the BS and ITMY (see his elog), I noticed that c1lsc was acting weird.  the IOP was the only one with the blinky heartbeat.  The IOP was all green lights, but all the other models had red for the fb connection, as well as the rightmost indicator (I don't know what that one is for).  I logged on to c1lsc and ran 'rtcds restart all'.  The script didn't get anywhere beyond saying it was beginning to stop the 1st model (sup, the bottom one on the lsc list).  Then all of the cpus went white.  I can still ping c1lsc, but I can't ssh to it.

I'm not sure what to do here Jamie.  Heelp. 

  7576   Thu Oct 18 15:36:57 2012 SteveUpdateCamerasWatec cameras & Tamron lenses

I purchased 3x  1/2" ccd cameras and 3x  F  50 mm lenses for the lab.

The spectral sensitivity plot is for an older model 902H. This new model has better sensitivity

Attachment 1: 10181201.PDF
10181201.PDF
  7575   Thu Oct 18 12:02:32 2012 DenUpdateIOOMCL, WFS triggers

Quote:

 

 I've redone the WFS triggers.  I left the MCL trigger alone (for now....I'll come back to it). 

The trigger was setup such that (a) it was totally unclear what was going on, by looking at the WFS screen.  Koji and I spent some time confused before I remembered that Den did this work recently.  Also, for some reason, the triggers were just plain thresholding, not a schmidt trigger, so any time the cavity flashed, the WFS came on.  Since the cavity can flash before the mcdown script has a chance to turn off the WFS servos, the outputs of the WFS filters are trying to output thousands of counts, and the signal goes through any time the cavity flashes.  Not so good.

 Your schmitt trigger has 2 threshold values - min and max. Set thresholding value in my trigger to the max of your schmitt trigger and you get the same behavior for MC,  triggers are not supposed to turn anything on in this realization as they do for locking with flashing.

  7574   Thu Oct 18 08:00:40 2012 jamieUpdateComputersRe: Lots of new White :(

Quote:

Solved. The power code of c1iscaux was loose.
Has anyone worked around the back side of 1Y3?


I looked into the problem. I went around the channel lists for each slow machines and found the variables are supported by c1iscaux

controls@pianosa:/cvs/cds/caltech/target/c1iscaux 0$ cd /cvs/cds/caltech/target/c1iscaux
controls@pianosa:/cvs/cds/caltech/target/c1iscaux 0$ grep C1:IF *
C1IFO_STATE.db:grecord(ai,"C1:IFO-STATE")

It seemed that the machine was not responding to ping. I went to 1Y3 and found the crate was off. Actually this is not correct.
The key was on but the power was off. I looked at the back and found the power code was loose from its inlet.
Once the code was pushed in and the crate was keyed, the white boxes got back online.

Just in case I burtrestored these slow channels by the snapshot at 6:07am on Sunday.

I was working around 1Y2 and 1Y3 when I wired the DAC in the c1lsc IO chassis in 1Y3 to the tip-tilt electronics in 1Y2.  I had to mess around in the back of 1Y3 to get it connected.  I obviously did not intend to touch anything else, but it's certainly possible that I did.

  7573   Thu Oct 18 03:57:20 2012 JenneUpdateLockingAlignment is really bad??

The goal of the night was to lock the Y arm.  (Since that didn't happen, I moved on to fixing the WFS since they were hurting the MC)

I used the power supplies at 1Y4 to steer PZT2, and watched the face of the black glass baffle at ETMY.  (elog 7569 has notes re: camera work earlier)  When I am nearly at the end of the PZT range (+140V on the analog power supply, which I think is yaw), I can see the beam spot near the edge of the baffle's aperture.  Unfortunately, lower voltages move the spot away from the aperture, so I can't find the spot on the other side of the aperture and center it.  Since the max voltage for the PZTs is +150, I don't want to go too much farther.  I can't take a capture since the only working CCD I found is the one which won't talk to the Sensoray.  We need some more cameras....they're already on Steve's list.

When the spot is a little closer to the center of the aperture than the edge of the aperture (so the full +150V!!), I don't see any beam coming out of AS....no beam out of the chamber at all, not just no beam on the camera.  Crapstick.  This is not good.  I'm not really sure how we (I?) screwed up this thoroughly.  Sigh.  Whatever ghost REFL beam that Kiwamu and Koji found last week is still coming out of REFL.

Previous PZT voltages, before tonight's steering:  +32V on analog power supply, +14.7 on digital.  This is the place that the PRMI has been aligned to the past week or so.

Next, just to see what happens, I think I might install a camera looking at the back (output) side of the Faraday so that I can steer PRM until the reflected beam is going back through the Faraday.  Team K&K did this with viewers and mirrors, so it'll be more convenient to just have a camera.

Advice welcome.

  7572   Thu Oct 18 03:45:56 2012 JenneUpdateIOOMCL, WFS triggers

Quote:

I've added MCL and WFS stop triggers into C1MCS/SUS model. Threshold value of MC_TRANS can be changed in the text entry located in MC2_POSITION medm screen. I tried 2 cases: trigger either blocks signal before MCL filter bank input or after output. Due to filter history in the 1 case MC2 was still slightly disturbed (C1:SUS-MC2_ULPD_VAR ~= 15) right after unlock. In the second case there was no disturbance as we zero output signal, but then I had to add "clear history" command to the mcup script.

WFS triggers block the signal before ASCPIT/YAW filter bank.

 I've redone the WFS triggers.  I left the MCL trigger alone (for now....I'll come back to it). 

The trigger was setup such that (a) it was totally unclear what was going on, by looking at the WFS screen.  Koji and I spent some time confused before I remembered that Den did this work recently.  Also, for some reason, the triggers were just plain thresholding, not a schmidt trigger, so any time the cavity flashed, the WFS came on.  Since the cavity can flash before the mcdown script has a chance to turn off the WFS servos, the outputs of the WFS filters are trying to output thousands of counts, and the signal goes through any time the cavity flashes.  Not so good.

I have removed the triggering for the angular DoFs from the mcs model (leaving the MCL triggering for now).  I have put new triggering into the ioo model, at the error point of the WFS loops.  The idea is that if the cavity unlocks, we don't want to lose the current pointing of the mirrors.  If the WFS servos were doing a lot of DC work, the bias sliders won't have the full information about where we want the mirrors to point.  Since we have the integrators in FM1, removing the input signal should freeze the output signal.  I need to modify the WFS on / off script so that this doesn't get turned off every lockloss.

Also, I have implemented (for the first time in a useful model, although I've done some testing in the tst model) the "wait" delay between a cavity locking and the trigger going through.  The idea is that we don't necessarily want the WFS to come on simultaneously with the cavity lock.  Since the wait delay resets any time it is un-triggered, this also prevents any signals from going through during cavity flashes.  The wait block has 3 inputs:  (1) a trigger, the output of some kind of trigger block, (2) a number of seconds to wait and (3) the model rate in Hz.  The model rate should be set with a constant in the model, the trigger passed from the trigger block, and the wait time in seconds should be available as an epics input. 

So far it looks like it's working as I expect, although I'm honestly too tired to do enough testing that I'm satisfied with, so I'm leaving the WFS off for the night.

  7571   Thu Oct 18 02:42:38 2012 DenUpdateSUS PITCH damping needed

Quote:

 

I think we can put ø2mm × 10mm long magnetic material inside 4 holes with actuation magnets. Then magnetic field on the other side of the mirror will be close to one produced by actuation magnet. Magnetic cylinder center of inertia will be in the vertical plane where mirror's center of inertia is. So this should not change alignment significantly. Eddy current dumping will be applied to the end of the magnetic cylinder opposite to the magnet using aluminium disks, we have them in the clean room.

 I've tested this approach. As we do not have required cylinders with high magnetic permittivity, I replaced them with magnets simular to actuator magnets ø2mm × 3mm long. Using them and aluminium disks from other TT I've made a "pitch dumping" construction.

DSC_4759.JPG

Pitch Q reduced but not that much as I could expect. I did a ringdown test. 

Plots:

yaw ringdown using original construction     |  yaw ringdown with added pitch damping

------------------------------------------------------------------------------------------------------------------------

pitch ringdown using original construction   |  pitch ringdown with added pitch damping

    yaw_nodamp_20.JPG   yaw_damped_13.JPG

pitch_nodamp_30.JPG     pitch_damped_10.JPG

 

 From this data I've estimated Q factor for yaw (135 vs 88) and pitch (192 vs 77) (original vs added pitch damping). Thess results diverges with the ones obtained by designes. They measured Q~40-50 for original construction. Pitch and yaw have 2 close resonances so this time domain method can not be very precise. I've measured the same with SR785.

yaw.pngpitch.png

In these comparison plots excitation was not the same as coils are not plugged in yet, but resonance Q factors can be compared.

 

  7570   Wed Oct 17 19:35:58 2012 KojiUpdateComputersRe: Lots of new White :(

Solved. The power code of c1iscaux was loose.
Has anyone worked around the back side of 1Y3?


I looked into the problem. I went around the channel lists for each slow machines and found the variables are supported by c1iscaux

controls@pianosa:/cvs/cds/caltech/target/c1iscaux 0$ cd /cvs/cds/caltech/target/c1iscaux
controls@pianosa:/cvs/cds/caltech/target/c1iscaux 0$ grep C1:IF *
C1IFO_STATE.db:grecord(ai,"C1:IFO-STATE")

It seemed that the machine was not responding to ping. I went to 1Y3 and found the crate was off. Actually this is not correct.
The key was on but the power was off. I looked at the back and found the power code was loose from its inlet.
Once the code was pushed in and the crate was keyed, the white boxes got back online.

Just in case I burtrestored these slow channels by the snapshot at 6:07am on Sunday.

  7569   Wed Oct 17 18:41:27 2012 JenneUpdateCamerasCamera looking at ETMY baffle

The camera titled "watec_mobile" is looking at the front of the black glass baffle (i.e. the side facing the ITM) on the ETMY table.  This required (for my quick hacky solution) removing the regular ETMYF camera.  Steve has a genius plan (I think) so that we can have both at the same time.  Anyhow, eventually we'll move the black glass back, so we'll be back to needing just one camera.

After dinner, I'll try aligning the Yarm.

  7568   Wed Oct 17 11:15:05 2012 DenUpdateSUS PITCH damping needed

Quote:

 

 Pitch damping solution needed! It should be in the machine shop already.

I think we can put ø2mm × 10mm long magnetic material inside 4 holes with actuation magnets. Then magnetic field on the other side of the mirror will be close to one produced by actuation magnet. Magnetic cylinder center of inertia will be in the vertical plane where mirror's center of inertia is. So this should not change alignment significantly. Eddy current dumping will be applied to the end of the magnetic cylinder opposite to the magnet using aluminium disks, we have them in the clean room.

pitch_damping.png

  7567   Wed Oct 17 09:01:37 2012 SteveUpdatePEMants on the PSL table

Quote:

We observed one or two ants climbing over PMC optics without booties and safety glasses.

The floor was mopped with strong Bayer Home Pest Control solution in the Vertex area.

Do not work inside the 40m lab if you are sensitive to chemicals!

 

 The floor is mopped again with strong PEST CONTROL SOLUTION in water in the Vertex area.

Do not plane to work in the IFO-room till noon if you are sensitive to chemicals!

  7566   Wed Oct 17 08:30:49 2012 SteveUpdateSUS PTICH damping needed

Quote:

Quote:

  [Koji, Steve, Den]

TT alignment is fine, yaw damping is satisfactory, pitch damping is slow. We might want to add magnets to the mirror and attach blades to the frame for pitch edge current damping.

We are moving towards electronics testing.

Atm1,  TT 1.5" high adaptor base will be back from the shop in 10 days.

Atm2,  There is no PITCH damping, YAW edie current damping works well at 0.5 mm gap

Atm3,  Adjustable Al -disc that contains a small magnet is purely designed.

We have to come up with a solution to have damping in PITCH

 Pitch damping solution needed! It should be in the machine shop already.

  7565   Wed Oct 17 08:05:10 2012 DenUpdateGeneraldo not leave op-table open

Quote:

ETMX optical table was left open. Burned toast award goes to ......?

 me

  7564   Wed Oct 17 08:04:31 2012 SteveUpdateVACRoughing pumps on for oil change

PR1, PR2 and RP3 turned on for warming up for oil change. Oil changed with 3.2L of  MVT-19 fluid in each. This substitute for HE-175 will be used at the next oil change - it has 1E-6 Torr vapor pressure.

To finish this job tomorrow: 1, check oil creeping upstream  2, change air filter of air purge if pressure drops <350 mTorr  3, measure venting time of pump

  7563   Wed Oct 17 08:01:54 2012 SteveUpdateGeneraldo not leave op-table open

ETMX optical table was left open. Burned toast award goes to ......?

  7562   Tue Oct 16 21:57:47 2012 DenUpdatePEMaccelerometers

All accelerometers are now at the table behind 1X4, cables are near readout box.

  7561   Tue Oct 16 20:40:06 2012 DenUpdateIOOc1lsc DAC0 now connected to tip-tilt SOS DW boards

  Full cable path from coil driver to osem input is now ready. I've tested Ch1-4 of the left AI and left coil driver. 15 pin outputs and monitors show voltage that we expect. I've checked voltage on the other side of the cable in the clean room, it is correct. We are ready to test the coils. We need to bake osem cables asap. Hopefully, Bob will start this job tomorrow.

DSC_4749.JPG DSC_4755.JPG

  7560   Tue Oct 16 17:13:23 2012 CzarinaSummaryGeneralvent stuff - 4 paths

I see 4+ possible paths for us to take, in terms of a possible vent in the next few weeks:

No Vent - Just do FPPRMI, using AS55

Mini Vent - Fix REFL path, nothing else.  ~1 day at atmosphere

Medium Vent - Fix REFL path, swap G&H mirrors for LaserOptik mirrors (so also resuspend passive TTs, maybe add pitch adjustment option). ~1 week or so at atmosphere - do this rather than Mini if Jan's Finesse calc says the G&H mirrors are too rough

Mega Vent - Fix all the things, do all the things.  Long time at atmosphere

The "+" is to take into account all the possible variations on "medium vent".  The No, Mini and Medium options assume we'll do the Mega option later, just not immediately.

  7559   Tue Oct 16 17:06:47 2012 SteveUpdateSAFETYsafety training

Ayaka Shoda, visiting graduate student received basic 40m specific safety training today

Attachment 1: IMG_1735.JPG
IMG_1735.JPG
  7558   Tue Oct 16 16:46:20 2012 Vent Czar and CzarinaSummaryGeneralvent stuff

We're discussing the plan for the next vent.  Now that we have all the tip-tilt stuff in hand, we should get ready for the big TT installation vent ASAP.  The question is what remains to be done, and when can we be ready

Stuff that needs doing:

* characterize TTs, check electrical connection of quadrupus cables, bake cables and base plates - a week or so (not including baking)

* phase map of LaserOptik mirrors (decide on needed resolution) - a week or so?

* get ready to swap G&H mirrors with LaserOptik mirrors (PR2, PR3, SR2, SR3)

* assembly of black glass beam dumps - one afternoon

* green periscope moving - day or so to make sure greens are resonant in arms, so we have something to align to

* redraw of REFL path?? (rotate 2nd refl mirror to reflect beam to the east, then add mirror to get beam out of IOO chamber viewport), cut hole in BS oplev table's box, install black beam tube

* FI camera mount - make sure it's here, baked

* active TT pitch damping plan

* passive TT weight addition and pitch alignment plan

* camera plan for taking in-chamber photos without touching the tables

* look at layout of ITMY table.  POY pick-off too close to main beam.  Can we move POY pick-off to BS table?

* remount black glass baffles on SOS cages.

 

Stuff to bake:

* Quadrupus cables

* TT base plates

* FI mirror mount for FI camera

* stuff to add to active TTs

* stuff to add to passive TTs

 

  7557   Tue Oct 16 11:54:05 2012 JenneUpdateIOOMC alignment??

The MC won't survive the boosts right now.  Pizza meeting is in a minute, and I won't be back to the lab before ~3:30 because of the seminar / a meeting, so someone else is welcome to try to fix it. Otherwise I'll have a look later on.

I'm leaving the autolocker disabled, so that it won't try any funny business.  WFS are off, so that they don't need to be turned off by the down script.

  7556   Tue Oct 16 11:38:17 2012 JenneUpdateLSCMore PRMI notes from last night

Quote:

How can you lock the PRMI without the REFL beams? c.f. this entry by Kiwamu
Which signals are you using for the locking?

I think the first priority is to find the fringes of the arms and lock them with POX/POY.

As for the POP, make sure the beam is not clipped because the in-vac steering mirrors
have been supposed to be too narrow to accommodate these two beams.

I was using AS55I for PRCL, and AS55Q for MICH.  I snuck that into the last line of an unrelated elog, since I did both things at the same time: see elog 7551.  Kiwamu's measurements (elog 6283) of the PRMI sensing matrix show that the PRCL and MICH signals are almost orthogonal in AS55 (although the optickle simulation doesn't agree with that...)  He was able to lock PRMI with AS55 I&Q (elog 6293), so I thought we should be able to as well.  Locking the PRMI was supposed to help tune the alignment of the PRM, not be the end goal of the night.  Also, we only tried locking PRCL in the "middle right" configuration, not the "lower left" configuration, but we were seeing what looked like recycling flashes only in the "lower left" configuration.

I agree in principle that we should be working on the arms. However, since we can't use the old steer-the-beam-onto-the-cage trick to find the beam, I was hoping that we could steer the beam around and see some light leaking out of the ETM, onto the end table.  However, with the 1% transmission of the ITMs and ~10ppm transmission of the ETMs, there's not a lot of light back there.  I was hoping to align the PRMI so that I get flashes with a gain of 10 if I'm lucky, rather than just the 5% transmission of the PRM.  With the PRMI aligned, I was expecting:

(1W  through Faraday) * (10 PR gain) * (0.5 BS transmission) * (0.01 ITM transmission) * (10ppm ETM transmission) = 0.5uW on the ETM tables during PRCL flashes

I was hoping that things would be well enough aligned that I could just go to the end table, and see the light with a viewer, although as I type this, I realize that if the beam was not on the end table (or even if it was...) any time I move the PZTs, I'd have to completely realign the PRMI in order to see the flashes.  This seems untenable, unless there are no other options.

We then got sidetracked by trying to see the POP beam, and once we saw the POP beam we wanted to put something down so we could find it again.  POP is also small, but not as small as expected at the end:

(1W  through Faraday) * (10 PR gain) * (20ppm PR2 transmission) = 0.2mW on POP during PRCL flashes.

POP was very difficult to see, and we were only able to see it by putting the foil in the beam path, and using a viewer.  I think that we once were able to see it by looking at a card with the viewer, but it's much easier with the foil.  I'd like to find an iris that is shiny (the regular black iris wasn't helpful), to facilitate this alignment.  Since we were just looking at the reflection off of the foil, I have no comment yet about clipping vs. not clipping.  I do think however that the forward-going beam may have been easier to find....when the PRMI alignment drifted, we lost the beam, but I could still see the forward-going beam.  Probably I should switch to that one, since that's the one that was lined up with the in-vac optics. 

Summary:

Ideas are welcome, for how to align the beam to the Yarm (and later to the Xarm), since our old techniques won't work.  Aligning the PRMI was a distraction, although in hopes of getting flashes so we could see some light at the end tables.  I'm going to go see if I can look through a viewport and see the edges of the black glass aperture, which will potentially be a replacement for the steering-on-the-cage technique, but if that doesn't work, I'm running out of ideas.

  7555   Tue Oct 16 02:34:38 2012 KojiUpdateLSCPOP lens placed on POX table, 2 PRMI movies

How can you lock the PRMI without the REFL beams? c.f. this entry by Kiwamu
Which signals are you using for the locking?

I think the first priority is to find the fringes of the arms and lock them with POX/POY.

As for the POP, make sure the beam is not clipped because the in-vac steering mirrors
have been supposed to be too narrow to accommodate these two beams.

  7554   Tue Oct 16 00:33:29 2012 JenneUpdateLSCPOP lens placed on POX table, 2 PRMI movies

[Evan, Jenne]

We aligned the PRMI.  We definitely can lock MICH, but we're not really sure if PRCL is really being locked or not.  I don't think it is.

Anyhow, we found 2 different places on the AS camera that we can align the PRMI.  One (middle, right hand side of the camera), we see the same weird fringing that we've been seeing for a week or two.  The other (lower left side of the camera), we see different fringing, almost reminds me more of back in the day a few months ago when the beam looked like it was expanding on each pass.  As I type, Evan is uploading the movies to youtube.  I *still* don't know how to embed youtube videos on the elog!

Also, we found both forward-going and backward-going POP beams coming out onto the POX table.  We placed the 2" lens in the path of the backwards beam, so that we can find it again.  We can't see it on an IR card, but if we put some foil where we think the beam should be, we can use a viewer to see the spot on the foil.  Poking a hole in the foil made an impromptu iris.

Youtube videos:

Lower left on camera

Middle right on camera

  7553   Tue Oct 16 00:08:26 2012 DenUpdateIOOc1lsc DAC0 now connected to tip-tilt SOS DW boards

Quote:

Tomorrow I'll get new SMA cables to connect the DW/AI outputs to the coil driver boards, and I'll start testing the coil driver outputs. 

 I've found a nice 16 twisted pair cable ~25m long and decided to use it as a port from 1Y3 to clean room cable instead of buying a new long one. I've added a break out board to the coil driver end to monitor outputs.

DSC_4748.JPG

  7552   Mon Oct 15 22:24:45 2012 JenneUpdateComputersLots of new White :(

Evan and I are starting to lock, and there is lots of new, unfortunate white stuff on several different screens.

C1:TIM-PACIFIC_STRING is gone, C1:IFO-STATE (MC state) is gone, C1:LSC-PZT..._requests are gone (all 4 of them), C1:PSL-FSS_FASTSWEEPTEST from the FSS screen is gone (although I'm not sure that that one is newly gone), lots of the WF AA lights on the LSC screen are gone.

Those are the things I find in a few minutes of not really looking around.

EDIT:  IPPOS is also gone, so I can't see how my current alignment relates to old alignments.

  7551   Mon Oct 15 22:16:09 2012 JenneUpdateSUSAll oplev gains turned to 0

Steve has promised to fix up all of the oplevs, but it hasn't happened yet, so I've turned all of the oplev gains to zero, so that when the optics are restored we don't have to quickly click them off.

Oplev values that were changed to zero:

PRM P=0.15, Y=-0.3

SRM P=-2.0, Y=2.0

BS P=0.2, Y=-0.2

ITMY P=2.1, Y=-2.0

ITMX P=1.0, Y=-0.5

ETMX P=-0.2, Y=-0.2

ETMY P=0.5, Y=0.6

Also, PRCL was changed in the LSC input matrix from REFL33I to AS55I, since there is no REFL beam out of the IFO :(

  7550   Mon Oct 15 20:45:58 2012 jamieUpdateIOOc1lsc DAC0 now connected to tip-tilt SOS DW boards

The tip-tile SOS dewhite/AI boards are now connected to the digital system.

20121015_190340.png

I put together a chassis for one of our space DAC -> IDC interface boards (maybe our last?).  A new SCSI cable now runs from DAC0 in the c1lsc IO chassis in 1Y3, to the DAC interface chassis in 1Y2.

Two homemade ribbon cables go directly from the IDC outputs of the interface chassis to the 66 pin connectors on the backplane of the Eurocrate.  They do not go through the cross-connects, cause cross-connects are stupid.  They go to directly to the lower connectors for slots 1 and 3, which are the slots for the SOS DW/AI boards.  I had to custom make these cables, or course, and it was only slightly tricky to get the correct pins to line up.  I should probably document the cable pin outs.

  • cable 0:  IDC0 on interface chassis (DAC channels 0-7) ---> Eurocrate slot 0 (TT1/TT2)
  • cable 1:  IDC1 on interface chassis (DAC channels 8-15)---> Eurocrate slot 2 (TT3/TT4)

As reported in a previous log in this thread, I added control logic to the c1ass front-end model for the tip-tilts.  I extended it to include TT_CONTROL (model part) for TT3 and TT4 as well, so we're now using all channels of DAC0 in c1lsc for TT control.

I tested all channels by stepping through values in EPICS and reading the monitor and SMA outputs of the DW/AI boards.  The channels all line up correctly.  A full 32k count output of a DAC channel results in 10V output of the DW/AI boards.  All channels checked out, with a full +-10V swing on their output with a full +-32k count swing of the DAC outputs.

   We're using SN 1 and 2 of the SOS DW/AI boards (seriously!)

The output channels look ok, and not too noisy.

Tomorrow I'll get new SMA cables to connect the DW/AI outputs to the coil driver boards, and I'll start testing the coil driver outputs.

As a reminder:  https://wiki-40m.ligo.caltech.edu/Suspensions/Tip_Tilts_IO

 

ELOG V3.1.3-